OpenStudy (anonymous):

Solve the equation 2x+1 = 30. Round your answer to the nearest hundredth. A. x ≈ 4.20 B. x ≈ 3.91 C. x ≈ 2.56 D. x ≈ 4.51

OpenStudy (anonymous):

14.5 isnt one of the answer choices thoug

OpenStudy (anonymous):

@gksriharsha

OpenStudy (blurbendy):

are you sure there's not a 14.5?

OpenStudy (anonymous):

yes I think the answer is A. thoug

OpenStudy (anonymous):

Do you mean: 2^(x + 1) = 30 x + 1 = log (base 2) of 30 x = (log (base2) of 30) - 1 x =~ 3.91

OpenStudy (blurbendy):

yes, that's right

OpenStudy (anonymous):

how about this one: Solve the equation 5x = 6.2. Round your answer to the nearest hundredth. A. x ≈ 1.13 B. x ≈ 2.59 C. x ≈ 1.45 D. x ≈ 3.49

OpenStudy (blurbendy):

5x^2 = 6.2 ?

OpenStudy (anonymous):

no 5x=6.2

OpenStudy (blurbendy):

well then x would be 1.24, but i don't see that

OpenStudy (anonymous):

round to the nearest hundreath, what would it be then?

OpenStudy (blurbendy):

1.24 still

OpenStudy (anonymous):

hmm well it should be one of those answers up their^^

OpenStudy (blurbendy):

well maybe it's like the last question where you forgot the ^2

OpenStudy (blurbendy):

do you mean 5^x = 6.2?

OpenStudy (blurbendy):

then the answer would be A

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

haha yay good e got that settled lol, how abou this one: Solve the equation of (¼)log6 (a – 3) – log6 3 = 0. A. 22 B. 64 C. 84 D. 96

OpenStudy (anonymous):

its 6^3 on last part of equation

OpenStudy (anonymous):

nevermind equation looks like this

OpenStudy (anonymous):